I already know how to get the section formula with the help of similarity of triangles 


concept, what i want to know know is whether i can get the same formula using the concept of distance between two coordinate points, 


i.e. for A(x1,y1) and  B(x2,y2) distance AB =     



 d=\sqrt{(\Delta x)^2+(\Delta y)^2}=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}.\,



now suppose that a point C divides AB in the ratio m : n, then using the distance formula 


   d=\sqrt{(\Delta x)^2+(\Delta y)^2}=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}.\, and using


  


can we get coordinates of C as




.

Grade:12

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Neel

You have to also apply the slope of AC= slope of CB

it will give you y-y1/x-x1 = y-y2/x-x2

as you are using only distance relation that is not sufficient.line ac and cb can be inclined also and can

 satisfy ac/cb=m/n

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..